OpenStudy (anonymous):

I don't understand how to do the first problem with the amount of commands provided by the first two lecture. I am not understanding how to solve this problem without using arrays

OpenStudy (anonymous):

What is the first problem?

OpenStudy (anonymous):

Python is really simple, don't over think it. You'll need to google search for details of commands, maybe grab a text book/ ebook or something. In C/C++ you'd need an array to do the first problem, in Python raw_input will just store the whole string. You'll also need to use the print command, real simple.

OpenStudy (anonymous):

I agree with Neil in that you shouldn't over think it. You should first create two string variables in which you will use the "raw_input" function in order to bind your input from the keyboard to your strings. There after use the print command to print out your two strings.

OpenStudy (e.mccormick):

The lectures assumed people knew a little more about programming. Remember, students in these classes had access to a 5 week class that was just on basic coding and taking in conjunction, a book, and recitation (smaller classes with a grad student leading them.) I suggest that people watch the first 10 to 15 videos here: https://buckysroom.org/videos.php?cat=36 They are short and cover individual topics. They are NOT rigorous, so they do not really cover the proper form of programming. But they are a good complement to getting you the very basics needed to make the class easier.

OpenStudy (anonymous):

I meant for the prime number problem sorry for not clarifying prime = [2] while len(prime) <= 1000: i=3 a = 0 for number in prime: testlist= [] testlist.append(i%number) if 0 in testlist: i=i+1 else: prime.append(i) i=i+1 print(prime[999]) for some odd reason this doesn't work. I can't seem to find an infinite loop though.

OpenStudy (e.mccormick):

Try with a small number, like looking for the first 25, and put in some print statments so you can see when different parts of the program are being called.